OpenStudy (anonymous):

NEED HELP 8TH GRADE MATH! A student is trying to solve the system of two equations given below: Equation P: y + z = 6 Equation Q: 3y + 4z = 1 Which of the following is a possible step used in eliminating the y-term? (y + z = 6) ⋅ 4 (3y + 4z = 1) ⋅ 4 (y + z = 6) ⋅ −3 (3y + 4z = 1) ⋅ 3

OpenStudy (anonymous):

@dogz_101 do you know this?

OpenStudy (anonymous):

I might im in 8th grade also so let me double look at it :)

OpenStudy (solomonzelman):

do you know what the elimination method is, have you ever done elimination previously?

OpenStudy (anonymous):

yes @SolomonZelman

OpenStudy (anonymous):

And thanks @dogz_101

OpenStudy (solomonzelman):

if so, then tell me - by what should you multiply Equation P, so that when you add Equation P and Equation Q you get one variable (lets say "y") eliminated?

OpenStudy (anonymous):

Im stuck on that thats why its hard i havent done any math in almost a week

OpenStudy (anonymous):

but i think its B, am I right? @SolomonZelman

OpenStudy (anonymous):

@dogz_101 did u find out?

OpenStudy (anonymous):

@kelliegirl33 do you know this?

OpenStudy (anonymous):

yes I do. you would multiply the first equation by a -3 this will give you : -3y - 3z = -18 3y + 4z = 1 --------------- 0 + z = -17 z = -17 as you can see, when added, the y terms cancel out...they are eliminated answer is C

OpenStudy (anonymous):

Thank you i have like 10 more of these this will help me alot THANK YOU!!

OpenStudy (anonymous):

sure thing :) tag me if you need help

OpenStudy (anonymous):

k thanks